An ultrasonic transducer of the type used in medical ultrasound imaging, is a very thin disk (m = 0.10 g) driven back and forth in SHM at 1.2 MHz by an electromagnetic coil. (a) The maximum restoring force that can be applied to the disk without breaking it is 36,000 N. What is the maximum oscillation amplitude that wont rupture the disk? (b) What is the disks maximum speed at this amplitude? X m/s
Answer #1

given

m = 0.1 g

f = 1.2 MHz

a. Fmax = 36,000 N

hence

maximum acceleration = amax = Fmax/m = 360000000 m/s/s

now, amax = Aw^2

here w = 2*pi*f

hence

A = 6.3325739*10^-6 m

b. maximum speed = Aw = 47.74648292 m/s